OpenStudy (anonymous):

The ratio of a quarterback's completed passes to attempted passes is 6 to 7. If he attempted 21 passes, find how many passes he completed. a. 7 passes c. 18 passes b. 25 passes d. 3 passes

Vocaloid (vocaloid):

set up a proportion completed passes/attempted passes = 6/7 = x/21 solve for x, which is the number of completed passes for 21 attempted passes

OpenStudy (anonymous):

how do i set it up

OpenStudy (mikeymaximum):

I think 18

OpenStudy (anonymous):

you are right

Vocaloid (vocaloid):

I just set up the proportion 6/7 = x/21

OpenStudy (anonymous):

Find each unit rate. $3.00 for 150 golf tees a. $0.02 per golf tee c. $50.00 per golf tee b. $153.00 per golf tee d. $0.2 per golf tee

Vocaloid (vocaloid):

price per tee = total cost/number of golf tees

OpenStudy (anonymous):

how would i set that up to solve it

OpenStudy (mikeymaximum):

I think that the answer is D

OpenStudy (mikeymaximum):

B and C are automatically out because they are higher than the price that they are giving us

OpenStudy (mikeymaximum):

No1 asked u... -_-

Vocaloid (vocaloid):

@joey1234567 what is the number of tees? what is the total cost?

OpenStudy (mikeymaximum):

When you are finding the unit rate then the answer that you are post 2 get should be smaller than the price that they are giving us

OpenStudy (mikeymaximum):

And Aa just doesn't make any sense at all because the price is just WAY to low

Vocaloid (vocaloid):

@joey1234567 if you still need help on this one unit price = cost/# of units = 3.00/150